OpenStudy (anonymous):

So the derivative is f'=4x+1 and it says to find the equation of the tangent line AND normal (perpendicular) line at x=2. so what exactly does it mean by the perpendicular line?

OpenStudy (mimi_x3):

m = 4(2) + 1 = 9 Tangent: y - y1 = m(x-x1) Normal: y - y1 = -1/m (x-x1) Do you mean perpendicular lines of the tangent to find the normal ?